ν_{1}+ν_{5} OF HCCN: DETERMINATION OF THE ν_{5} VIBRATIONAL ENERGY

\underline{JIAXIANG HAN}, P. Y. HUNG, JOHN DeSAIN, W. E. JONES, and R. F. CURL, Chemistry Department and Rice Quantum Institute, Rice University, Houston, Texas 77251.

The high resolution infrared spectrum of HCCN in the region 3338- 3382cm^{−1} has been observed by infrared kinetic spectroscopy. The spectrum is assigned as the combination band ν_{1}+ν_{5} of the quasilinear molecule HCCN with the origin at 3355.510(9)cm^{−1}. Based on information from the ν_{1}+ν_{5}-ν_{5} spectrum^{[1]}, the energy of the lowest excited state with angular momentum about the a-axis, ν_{5}, is determined to be 128.913(9)cm^{−1}. This value is lower than the results obtained by means of relative intensity measurements on the millimeter-wave spectra^{[2]} [145(15)cm^{−1}] or from similar relative intensity measurements on the IR spectra^{[1]} [187(20)cm^{−1}]. The present value of the energy for ν_{5} predicts a barrier to linearity similar to that obtained from the corresponding band of DCCN^{[3]} and is higher than that found by McCarthy et al.^{[2]}.

[1] C. L. Morter, S. K. Farhat and R. F. Curl, Chem. Phys. Lett. 207, 153(1993); ^{[2]} M. C. McCarthy, C. A. Gottlieb, A. L. Cooksy and P. Thadeus, J. Chem. Phys. 103(18), 7779(1995); ^{[3]} Fuge Sun, A. Kosterev, G. Scott, V. Litosh, and R. F. Curl, J. Chem. Phys. 109(20), 8851(1998).